# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v3

@@ 4.3 Comparison of Computational Time between Constant BS and Increasing BS via SFO Complexity
Figure 5: Norm of objective function gradient versus SFO complexity. Datasets used were COIL100 (PCA), MNIST (PCA), MovieLens-1M (LRMC), and Jester (LRMC) in order from left to right. A cosine annealing LR was used except COIL100, for which a constant LR was used. For ‘BS-increases=0=0,’ a constant BS $b=b_{0}$ was used. For ‘BS-increases=3=3’ and ‘BS-increases=6=6,’ the BS was increased 33 and 66 times, respectively, in accordance with the exponential-growth BS schedule.
@@
Large BS 1411.37 2697.24 4482.41 5432.54
Full BS 1488.34 2719.66 12947.91 6816.28
@@
Table 2: Computational time [sec] (CPU time) for large, full, and increasing batch size.
@@
What are the differences among using a large constant BS, a small constant BS, and an increasing BS? We numerically investigate this question on the basis of SFO complexity, which is defined by the number of stochastic gradient evaluations executed over $T$ iterations Agarwal & Bottou (2015); Shallue et al. (2019); Sato & Iiduka (2023). For a constant BS $b$, it is represented by $bT$. SFO complexity for other BSs can be calculated numerically, serving as a proxy for computational time. Figure 5 plots the gradient norm of the objective function against SFO complexity for PCA and LRMC. Each curve corresponds to RSGD for $3000$ steps under one of the following settings: constant BS, BS tripled every $1000$ steps (three increases in total), or BS tripled every $500$ steps (six increases in total). These settings follow the update formula for constant or exponential growth BS. Additional experimental details are given in the caption of Figure 5. As shown in the figure, increasing BS combines the advantages of both small and large BSs, namely shorter computational time and convergence to a solution with a smaller final gradient norm).
@@
As shown in Figure 5, although both full and large constant BSs lead to minimizers achieving a small gradient norm, they require more SFO complexity. In contrast, while small constant BSs require fewer SFO complexity, they do not lead to minimizers achieving a small gradient norm. Increasing BSs with many increments and a small initial batch size lead to minimizers achieving a small gradient norm with relatively few SFO complexity. As shown in Table 2, the computational time (CPU time) of increasing BS is shorter than that of both large and full constant BSs. Between small and large constant BSs, there is a trade-off between convergence to a solution with a smaller final gradient norm and a shorter computational time. Our results demonstrate that an increasing BS balances this trade-off effectively.
@@
Why is increasing BS better?  Our theoretical analyses provide an answer to this question. From Theorem 1, the SFO complexity of a constant BS with the critical BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\epsilon^{-4})$. From the same theorem, the SFO complexity of an exponential growth BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\gamma^{\epsilon^{-2}})$, where $\gamma\coloneq 3$ in this experiment. Although $O(\gamma^{\epsilon^{-2}})$ is inferior to $O(\epsilon^{-4})$ in our strict theoretical setting—fixed $K$ (number of steps for each BS) and dynamic $M$ (number of BS increases)—the assumptions in this experiment differed. Specifically, $K$ depended on the value assigned to $M$, and our analysis can be applied to this experimental setting. Under these conditions, the SFO complexity of an exponential growth BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\epsilon^{-2})$, which is superior to $O(\epsilon^{-4})$. This provides our theoretical justification: increasing BS achieves lower SFO complexity than constant BS. Moreover, even among increasing BS settings, performance was better for $M=6$ than for $M=3$; i.e., performance improves with the number of BS increments. From Theorem 2, we have $\|G_{T}\|=O(\sqrt{\frac{1}{T}+\frac{1}{M}})$, indicating that a larger $M$ leads to a smaller gradient norm and explains the observed phenomenon. The relevant derivations are provided in Appendix A.
Remark 2.
@@
Under the rigorous conditions of our theory, $M\to\infty$ implies that $T=MK+\text{const}.\to\infty$ with $K$ fixed. In contrast, the conditions of our experiment do not permit an increase in BS to achieve $T\to\infty$ because it requires $K\to\infty$ for $T$ to diverge with a fixed $M$, which means that the number of steps for the initial batch size is infinite. However, as is done in many experimental setups, including ours, $T$ is predetermined as a finite value before conducting the experiment. This setting is equivalent to using the number of steps or epochs as the stopping criterion. Under this practical setting, the conditions of our experiment are meaningful because $T<\infty$ implies $K<\infty$ with a fixed $M$ (therefore, $K$ is a fixed finite value). For these reasons and from our proofs under a practical setting, our theoretical analysis can be used to evaluate the performance of RSGD including calculation of SFO rates under our experimental setting. Thus, our theoretical analysis supports the numerical results.
@@ 5 Conclusion
Our theoretical analysis using several learning rates, including cosine annealing and polynomial decay, demonstrated that using an increasing batch size rather than a constant batch size improves the convergence rate of Riemannian stochastic gradient descent. This finding is supported by our experimental results. Furthermore, increasing batch size yields minimizers that achieve a smaller gradient norm within a shorter computational time owing to improvements in the stochastic first-order oracle rate. These results indicate that increasing batch size combines the advantages of both small and large batch size. However, due to the nature of the experimental tasks, we were unable to investigate its effect on generalization performance. We believe our results, which clarify one facet of the effectiveness of increasing batch size, offer insight into this important problem.
@@ Appendix A Calculation of SFO Complexity
A.1 SFO Rates with Constant BS and Increasing BS
We compute the SFO rate required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}(\epsilon>0)$, as described in Section 4.3.
@@ A.1 SFO Rates with Constant BS and Increasing BS
From Theorem 1, the number of iterations $T$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is
$T=\frac{\tilde{Q}_{1}}{\epsilon^{2}-\tilde{Q}_{2}\sigma b^{-1}}.$
Since the SFO complexity of constant BS can be represented by $\text{SFO}(b)=bT$, the $\text{SFO}(b)$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is given by
$\text{SFO}_{\epsilon}(b)=bT=b\frac{\tilde{Q}_{1}}{\epsilon^{2}-\tilde{Q}_{2}\sigma b^{-1}}=\frac{b^{2}\tilde{Q}_{1}}{b\epsilon^{2}-\tilde{Q}_{2}\sigma}.$
Since
$\frac{d}{db}\text{SFO}_{\epsilon}(b)=\frac{2b\tilde{Q}_{1}(b\epsilon^{2}-\tilde{Q}_{2}\sigma)-b^{2}\tilde{Q}_{1}\epsilon^{2}}{(b\epsilon^{2}-\tilde{Q}_{2}\sigma)^{2}}=\frac{b\tilde{Q}_{1}(b\epsilon^{2}-2\tilde{Q}_{2}\sigma^{2})}{(b\epsilon^{2}-\tilde{Q}_{2}\sigma)^{2}}$
@@
$b\leq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}\Rightarrow\frac{d}{db}\text{SFO}_{\epsilon}(b)\leq 0\quad\text{and}\quad b\geq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}\Rightarrow\frac{d}{db}\text{SFO}_{\epsilon}(b)\geq 0,$
we have the critical batch size of constant BS:
$b^{\star}\coloneq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}=\underset{b>0}{\text{argmin}}\text{SFO}_{\epsilon}(b),$
@@
$\text{SFO}_{\epsilon}(b^{\star})=\frac{4\tilde{Q}_{1}\tilde{Q}_{2}^{2}\sigma^{4}\epsilon^{-4}}{\tilde{Q}_{2}\sigma^{2}}=4\tilde{Q}_{1}\tilde{Q}_{2}\sigma^{2}\epsilon^{-4}=O(\epsilon^{-4}).$
@@
From Theorem 1, the number of iterations $T$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is given by
@@
We assume a setting in which $T=MK$, meaning that by the $T$-th iteration, the BS has been increased $M$ times, and RSGD has been updated for $K$ steps at each BS. If the exact value of $T$ takes the form $T=MK+k(1\leq k\leq K-1)$, we can set $T$ to $M(K+1)$ to ensure that $\|G_{T}\|^{2}\leq\epsilon^{2}$ still holds. Similarly, if $T$ takes the form $T=MK\textendash k(1\leq k\leq K-1)$, we can set $T$ to $MK$ to satisfy the same condition. Hence, our assumption that $T=MK$ is reasonable.
@@
$\text{SFO}=\left(\sum_{m=0}^{M-1}b_{0}\gamma^{m}\right)K=\frac{b_{0}}{\gamma-1}K(\gamma^{M}-1).$
@@
$\text{SFO}_{\epsilon}=\frac{b_{0}K}{\gamma-1}(\gamma^{\frac{\tilde{Q}_{1}+\tilde{Q}_{2}\sigma^{2}b_{0}^{-1}}{K\epsilon^{2}}}-1)=O(\gamma^{\epsilon^{-2}}).$
Since $\tilde{Q}_{2}$ contains $K$, as shown by the formal statement of Theorem 2 (with constant LR and exponential growth BS), we rearrange $\tilde{Q}_{2}$ for calculating SFO rate under our experimental setting as
@@
Therefore, under our experiment setting—where $K$ is determined by fixed MM—and with $K=\frac{T}{M}$, the SFO complexity of increasing BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ becomes
$\text{SFO}_{\epsilon}=\frac{b_{0}(\gamma^{M}-1)}{\gamma-1}\frac{\tilde{Q}_{1}}{M\epsilon^{2}-\tilde{Q}_{3}\sigma^{2}b_{0}^{-1}}=O(\epsilon^{-2}).$
A.2 Relationship between $M$ the number of increments the BS and the convergence of RSGD
In our numerical experiment in Section 4.3, among the increasing BS configurations, the one with six total BS increases exhibited notably better performance than the one with three increases. We explain this result on the basis of our theoretical analysis. In this section, we assume the same condition as in A.1; namely, $T=MK$. From Theorem 1, we have the following bound for constant and exponential growth BS:
$\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}]$ $\leq\frac{1}{(2-L_{r}\eta_{\max})\eta_{\max}}\left\{2(f(x_{0})-f^{\star})+\frac{L_{r}\eta_{\max}^{2}K\gamma}{\gamma-1}\frac{\sigma^{2}}{b_{0}}\right\}\frac{1}{T}$
$=\frac{\tilde{Q}_{1}}{T}+\tilde{Q}_{3}\sigma^{2}b_{0}^{-1}\frac{K}{T}=\frac{\tilde{Q}_{1}}{T}+\tilde{Q}_{3}\sigma^{2}b_{0}^{-1}\frac{1}{M}$
$=O\left(\frac{1}{T}+\frac{1}{M}\right).$
This rate suggests larger $M$ leads to faster convergence rate. From this perspective, it is reasonable that $M=6$ outperformed $M=3$ in our numerical experiment.
Appendix B Objective Function Value against SFO Complexity

### v4

@@ 4.3 Comparison of Computational Time between Constant BS and Increasing BS versus SFO Complexity
Figure 5: Norm of objective function gradient versus SFO complexity. Datasets used were COIL100 (PCA), MNIST (PCA), MovieLens-1M (LRMC), and Jester (LRMC) in order from left to right. A cosine annealing LR was used except for COIL100, for which a constant LR was used. For ‘BS_increases=0=0,’ a constant BS $b=b_{0}$ was used. For ‘BS_increases=3=3’ and ‘BS_increases=6=6,’ the BS was increased 33 and 66 times, respectively, in accordance with the exponential growth BS.
@@
Large constant BS 1411.37 2697.24 4482.41 5432.54
Full constant BS 1488.34 2719.66 12947.91 6816.28
@@
Table 2: Computational time [s] (CPU time) for large-constant, full-constant and increasing batch sizes.
@@
What are the differences between using a large constant BS, a small constant BS, and an increasing BS? We numerically investigated this question on the basis of SFO complexity, defined as the number of stochastic gradient evaluations executed over $T$ iterations (Agarwal and Bottou, 2015; Shallue et al., 2019; Sato and Iiduka, 2023). For a constant BS $b$, SFO complexity is represented by $bT$. For other BSs, it can be computed numerically, serves as a proxy for computational time. Figure 5 plots the gradient norm of the objective function versus SFO complexity for PCA and LRMC. Each curve corresponds to RSGD for $3000$ steps under one of the following settings: constant BS, BS tripled every $1000$ steps (three increases in total), or BS tripled every $500$ steps (six increases in total). These settings follow the update formula for constant or exponential growth BS. Additional experimental details are provided in the caption of Figure 5. As shown in the figure, an increasing BS combines the advantages of both small and large BSs—namely shorter computational time and convergence to a solution with a smaller final gradient norm.
@@
As shown in Figure 5, although both full and large constant BSs lead to optima achieving a small gradient norm, they require more SFO complexity. In contrast, while small constant BSs require less SFO complexity, they do not lead to optima achieving a small gradient norm. As shown in Table 2, the computational time (CPU time) of an increasing BS is shorter than that of both large and full constant BSs. Between small and large constant BSs, there is a trade-off between convergence to a solution with a smaller final gradient norm and a shorter computational time. Our results show that an increasing BS balances this trade-off effectively.
@@
Why is increasing BS better? Our theoretical analyses provide an answer to this question. From Theorem 3.2, the SFO complexity of a constant BS with the critical BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\epsilon^{-4})$. From Theorem 3.3, the SFO complexity of an exponential growth BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\gamma^{\epsilon^{-2}})$, where $\gamma\coloneq 3$ in this experiment. Although $O(\gamma^{\epsilon^{-2}})$ is inferior to $O(\epsilon^{-4})$ in our strict theoretical setting—with fixed $K$ (number of steps for each BS) and dynamic $M$ (number of BS increases)—the assumptions in this experiment differ. Specifically, $K$ depended on the value assigned to $M$, and our analysis can be applied to this experimental setting. Under these conditions, the SFO complexity of an exponential growth BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ is $O(\epsilon^{-2})$, which is superior to $O(\epsilon^{-4})$ and consistent with our experimental results. This provides our theoretical justification: increasing BS achieves lower SFO complexity than a constant BS. Summarizing the above, we obtain the following theorem. The derivation is provided in Appendix A.
Theorem 4.1.
Let $\epsilon>0$, and consider an LR scheduler other than the diminishing LR one. Under the assumptions of Theorems 3.2 and 3.3, the SFO complexity for achieving $\|G_{T}\|^{2}<\epsilon^{2}$ is $O(\epsilon^{-4})$ for a constant BS and $O(\epsilon^{-2})$ for an increasing BS.
Remark 4.2.
@@
Under the rigorous conditions of our theory, $M\to\infty$ implies that $T=MK+\text{const}.\to\infty$ with fixed $K$. In contrast, the conditions of our experiment do not permit an increase in BS to achieve $T\to\infty$ because it would require $K\to\infty$ for $T$ to diverge with fixed $M$, which means that the number of steps for the initial BS is infinite. However, as is done in many experimental setups, including ours, $T$ is predetermined as a finite value before conducting the experiment. This setting is equivalent to using the number of steps as the stopping criterion. Under this practical setting, the conditions of our experiment are meaningful because $T<\infty$ implies $K<\infty$ with a fixed $M$. On the other hand, by considering the conditions of our proofs, our theoretical analysis can also be applied to calculate SFO complexities, even within our experimental setting. From these reasons, our theoretical analysis can support the numerical results. To enhance understanding, the derivation of the SFO complexity for an increasing BS can be found in Appendix A.
Figure 6: The plot on the far left depicts the objective function value (loss) for MNIST (PCA) under the same setting as in Figure 5. The plot in the middle left (resp. right) depicts the objective function value (loss) for our proposed optimization problem when $(N,n)=(7200,1024)$ (resp. $(60000,1024)$). The plot on the far right depicts the trade-off with respect to the setting of $\gamma$.
$(b_{0},M)$ (27,0) (243, 0) (6561, 0) (7200, 0) (27, 3) (729, 3) (800, 3) (27, 6) (30, 6)
$x$ (COIL100) 2.233 1.055 0.903 0.902 1.173 0.911 0.902 0.973 0.951
$(b_{0},M)$ (243, 0) (2187, 0) (59049, 0) (60000, 0) (243, 3) (2223, 3) (6561, 3) (243, 6) (247, 6)
$y$ (MNIST) 4.846 2.292 4.847 5.332 1.141 -4.622 2.134 4.247 -4.139
Table 3: Let $U_{\text{EVD}}$ be the solution to PCA by eigenvalue decomposition, and let $U_{\text{RSGD}}$ be the solution to PCA by RSGD. Then, the differences between the representations of the subspaces spanned by the principal components of $U_{\text{EVD}}$ and $U_{\text{RSGD}}$ are given by $l\coloneq\|U_{\text{EVD}}U_{\text{EVD}}^{\top}-U_{\text{RSGD}}U_{\text{RSGD}}^{\top}\|$. For comparison with the results for COIL100, we introduce $x$ such that $l=9.487+x\times 10^{-3}$, and for MNIST, we introduce $y$ such that $l=0.4+y\times 10^{-4}$. $M$ is the number of BS increments.
Objective Function Values: Generalization performance is an important issue, and the objective function value serves as a criterion for measuring it. As shown in part of Appendix A.2, although the exponential growth BS performed better than or equal to the constant BS with respect to the objective function values, the differences in the values were small. One possible reason for this is that the objective function may be flat around the optimal solution. To test this hypothesis, we conducted an experiment comparing the solution obtained by RSGD with that obtained by eigenvalue decomposition in PCA. When the data dimensionality is not large, PCA can obtain the exact solution by eigenvalue decomposition, and the difference from the approximate solution obtained by RSGD can be used as a measure of generalization. As shown in Table 3, the difference is small, which indicates that the objective functions of PCA with COIL100 and MNIST are flat around the optimal (exact) solution. Conversely, by considering optimization problems in which the objective functions do not have this property, the superiority of an increasing BS in terms of generalization performance can be revealed more clearly. In fact, this has already been demonstrated in LRMC with MovieLens-1M, as shown at the far left of Figure 6 (the corresponding results for the other datasets are provided in Appendix A.2). For further verification of this, we formulated a new optimization problem on the sphere $\text{S}^{n-1}$: $\text{minimize}_{w\in\text{S}^{n-1}}f(w)\coloneq\frac{1}{N}\sum_{j=1}^{N}\sqrt{|\langle x_{j},w\rangle|}$, where $\{x_{j}\}_{j=1}^{N}$ is a dataset uniformly sampled from $\text{S}^{n-1}$ (see Appendix C for more details). We set $(N,n)=(7200,1024)$ and $(60000,1024)$ and used the cosine annealing LR with an initial LR $\eta_{\max}=0.01$. The BS was configured in exactly the same way as that for the COIL100 and MNIST datasets in Section 4.3. As in the two central graphs of Figure 6, an increasing BS achieved superior performance in terms of the objective function value for this proposed problem. Although the order of the differences (vertical axis) $O(10^{-2})-O(10^{-1})$ may appear small, it is 10–100 times larger than that observed for PCA because the differences in the values shown in Table 3 are on the order of $O(10^{-4})-O(10^{-3})$. These results further confirm the validity of our hypothesis for this problem. Given these results, we suggest that an increasing BS should also improve generalization performance. Turning to a different topic, the objective function of our proposed optimization problem has an unbounded gradient norm (see Appendix C), and the corresponding experiments numerically confirmed that our theoretical results are applicable even in this case. The graphs of the gradient norm for our proposed problem are provided in Appendix C.1.
Guidelines for setting $\gamma,b_{0}$ and $M$ in the exponential growth BS : From Theorem 3.3, $\|G_{T}\|^{2}=O(1+\frac{\gamma}{\gamma-1})=O(1+b_{0}^{-1})=O(M^{-1})$ and $\text{SFO}^{\text{incr}}_{\epsilon}=O(\gamma^{M})=O((b_{0}^{2}-1)^{-1}b_{0}^{3})=O(\gamma^{M}M^{-1})$ hold. These theoretical results suggest a trade-off: for smaller gradient norm, $\gamma,b_{0},$ and $M$ should be assigned larger values while for smaller SFO complexity, $\gamma,b_{0}$, and $M$ should be assigned smaller values. The far-right plot in Figure 6 illustrates the trade-off between the gradient norm and SFO complexity achieved experimentally on MNIST (PCA) using the cosine annealing LR with an initial LR $\eta_{\max}=0.01$, where $(b_{0},M)=(247,6)$, while varying $\gamma$ among $\{2.0,2.5,3.0,3.5,4.0\}$. For $b_{0}$, this trade-off can be observed by comparing the three entries under ‘BS_increases=3=3’ or the two entries under ‘BS_increases=6=6’ in Figure 5. For $M$, which represents the value ‘BS_increases’, this trade-off can be observed by comparing the purple entry with the pink entry in Figure 5. From these results, we were able to show, both theoretically and experimentally, that for each hyperparameter the trade-off between the gradient norm and SFO complexity holds. Figure 5 shows that an increasing BS with small $b_{0}$ and large $M$ $(=6)$ leads to multiple optima characterized by a small gradient norm and relatively low SFO complexity, indicating that well-balanced configurations were achieved by taking into account the trade-offs inherent in $b_{0}$ and $M$ individually. The relevant derivation is provided in Appendix B.
@@ 5 Conclusion
Our theoretical analysis, conducted with several learning rate schedules, including cosine annealing and polynomial decay, demonstrated that using an increasing batch size rather than a constant batch size improves the convergence rate of Riemannian stochastic gradient descent. This result is supported by our experimental results. Furthermore, an increasing batch size yields optima characterized by a smaller gradient norm within a shorter computational time owing to reductions in the stochastic first-order oracle complexity. These findings indicate that an increasing batch size combines the advantages of both small and large constant batch size. Due to the nature of the experimental tasks, we were unable to directly investigate the effect on generalization performance. However, evaluation via the objective function value suggests that an increasing batch size also enhances generalization performance. We believe our results, which clarify one aspect of the effectiveness of an increasing batch size for generalization performance, provide valuable insight into this topic.
[…]
This work was supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number 24K14846 awarded to Hideaki Iiduka.
@@ Appendix A Calculation of SFO Complexity
A.1 SFO Complexity with Constant BS and Increasing BS
We compute the SFO complexity required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ $(\epsilon>0)$, as described in Section 4.3.
@@ A.1 SFO Complexity with Constant BS and Increasing BS
From Theorem 3.2, the number of iterations $T$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is
$T=\frac{\tilde{Q}_{1}}{\epsilon^{2}-\tilde{Q}_{2}\sigma^{2}b^{-1}}.$
Because the SFO complexity of constant BS can be represented by $\text{SFO}^{\text{const}}(b)=bT$, the $\text{SFO}^{\text{const}}(b)$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is given by
$\text{SFO}_{\epsilon}^{\text{const}}(b)=bT=b\frac{\tilde{Q}_{1}}{\epsilon^{2}-\tilde{Q}_{2}\sigma^{2}b^{-1}}=\frac{b^{2}\tilde{Q}_{1}}{b\epsilon^{2}-\tilde{Q}_{2}\sigma^{2}}.$
Because
$\frac{d}{db}\text{SFO}_{\epsilon}^{\text{const}}(b)=\frac{2b\tilde{Q}_{1}(b\epsilon^{2}-\tilde{Q}_{2}\sigma^{2})-b^{2}\tilde{Q}_{1}\epsilon^{2}}{(b\epsilon^{2}-\tilde{Q}_{2}\sigma^{2})^{2}}=\frac{b\tilde{Q}_{1}(b\epsilon^{2}-2\tilde{Q}_{2}\sigma^{2})}{(b\epsilon^{2}-\tilde{Q}_{2}\sigma^{2})^{2}}$
@@
$b\leq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}\Rightarrow\frac{d}{db}\text{SFO}_{\epsilon}^{\text{const}}(b)\leq 0\quad\text{and}\quad b\geq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}\Rightarrow\frac{d}{db}\text{SFO}_{\epsilon}^{\text{const}}(b)\geq 0,$
we have the critical BS of constant BS:
$b^{\star}\coloneq\frac{2\tilde{Q}_{2}\sigma^{2}}{\epsilon^{2}}=\underset{b>0}{\text{argmin}}\ \text{SFO}_{\epsilon}^{\text{const}}(b),$
@@
$\text{SFO}_{\epsilon}^{\text{const}}(b^{\star})=\frac{4\tilde{Q}_{1}\tilde{Q}_{2}^{2}\sigma^{4}\epsilon^{-4}}{\tilde{Q}_{2}\sigma^{2}}=4\tilde{Q}_{1}\tilde{Q}_{2}\sigma^{2}\epsilon^{-4}=O(\epsilon^{-4}).$
@@
From Theorem 3.3, the number of iterations $T$ required to achieve $\|G_{T}\|^{2}\leq\epsilon^{2}$ is given by
@@
We assume a setting in which $T=MK$, meaning that by the $T$-th iteration, the BS has been increased $M$ times, and RSGD has been updated for $K$ steps at each BS. If the exact value of $T$ takes the form $T=MK+k$ $(1\leq k\leq K-1)$, we can set $T$ to $M(K+1)$ to ensure that $\|G_{T}\|^{2}\leq\epsilon^{2}$ still holds. Similarly, if $T$ takes the form T=MKkT=MK–k $(1\leq k\leq K-1)$, we can set $T$ to $MK$ to satisfy the same condition. Hence, our assumption that $T=MK$ is reasonable.
@@
$\text{SFO}^{\text{incr}}=\left(\sum_{m=0}^{M-1}b_{0}\gamma^{m}\right)K=\frac{b_{0}}{\gamma-1}K(\gamma^{M}-1).$
@@
$\text{SFO}_{\epsilon}^{\text{incr}}=\frac{b_{0}K}{\gamma-1}(\gamma^{\frac{\tilde{Q}_{1}+\tilde{Q}_{2}\sigma^{2}b_{0}^{-1}}{K\epsilon^{2}}}-1)=O(\gamma^{\epsilon^{-2}}).$
Because $\tilde{Q}_{2}$ contains $K$, as shown by the formal statement of Theorem 3.3 (with a constant LR and an exponential growth BS), we rearrange $\tilde{Q}_{2}$ for calculating SFO complexity under our experimental setting as
@@
Therefore, under our experimental setting—where $K$ is determined by fixed MM—and with $K=\frac{T}{M}$, the SFO complexity of increasing BS for achieving $\|G_{T}\|^{2}\leq\epsilon^{2}$ becomes
$\text{SFO}_{\epsilon}^{\text{incr}}=\frac{b_{0}(\gamma^{M}-1)}{\gamma-1}\frac{\tilde{Q}_{1}}{M\epsilon^{2}-\tilde{Q}_{3}\sigma^{2}b_{0}^{-1}}=O(\epsilon^{-2}).$
The calculations in this section are for the case of a constant LR, but it is immediately clear from Theorems 3.2 and 3.3 that exactly the same results hold for a cosine annealing LR, which was also used in our experiments.
A.2 Objective Function Value versus SFO Complexity